find the values of k for which the following equations have two equal roots.

(a)4x^2+4(k-2)x+k=0

(b)(k+2)x^2+4k=(4k+2x)

(c)x^2-2x+1=2k(k-2)

(d)(k+1)x^2+kx-2k=0

(e)4x^2-(k-2)x+9=0


1
Expert's answer
2022-01-09T16:59:53-0500

(a)


"4x^2+4(k-2)x+k=0"

"D=(4(k-2))^2-4(4)(k)=0"

"k^2-4k+4-k=0"

"k^2-5k+4=0"

"(k-1)(k-4)=0"

"k\in \{1, 4\}"

(b)


"(k+2)x^2+4k=4k+2x"

"k\not=-2"


"(k+2)x^2-2x=0"

"x_1=0, x_2=\dfrac{2}{k+2}\not=0, k\in \R, k\not=-2"

There is no solution.


"k\in \empty"

(c)


"x^2-2x+1=2k(k-2)"

"(x-1)^2=2k(k-2)"

"2k(k-2)=0"

"k_1=0, k_2=2"

"k\in \{0,2\}"

(d)


"(k+1)x^2+kx-2k=0"

"k\not=-1"


"D=(k)^2-4(k+1)(-2k)=0"

"k^2+8k+8=0"

"k^2+8k+16=8"

"(k+4)^2=8"

"k_1=-2\sqrt{2}, k_2=2\sqrt{2}"

"k\in \{-2\sqrt{2},2\sqrt{2}\}"

(e)


"4x^2-(k-2)x+9=0"

"D=(-(k-2))^2-4(4)(9)=0"

"(k-2)^2=144"

"k_1=-10, k_2=14"

"k\in \{-10,14\}"